OpenStudy (anonymous):

Austin wants to sketch the graph of a circle represented by the given equation. (x - 4)2 + (y + 7)2 = 64 Using complete sentences, describe the steps used to sketch the circle on a coordinate grid.

OpenStudy (anonymous):

according to your given... do you know where Center is and what dimension radius has? :)

OpenStudy (anonymous):

no how do u find that out agein?

OpenStudy (anonymous):

i really have no idea how to do this i dont understand it at all

OpenStudy (anonymous):

\[\LARGE (x-h)^2+(y-k)^2=r^2\] Center (h,k) Radius (r)

OpenStudy (anonymous):

radius would be 64 and center would be (4,7) then right?

OpenStudy (anonymous):

or would the senter be (-4,7)?

OpenStudy (anonymous):

not actually... in your case you have: \[\LARGE (x-4)^2+[y-(-7)]^2=8^2\] so Center is 4,7 and radius 8

OpenStudy (anonymous):

oh alright

OpenStudy (anonymous):

sorry Center is 4,-4 ... and radius 8 :)

OpenStudy (anonymous):

AHAHAHAHAHAHAHAAHAHHAA...... sorry Center is 4,-7 ... and radius 8 :) I'm in hurry... typo mistake .. lol

OpenStudy (anonymous):

all good lol

OpenStudy (anonymous):

ok .. now let's start sketching... we know that center is C(4.-7) which means x=4 and y=-7 ... and this looks like:  so center of the circle is at point P.